# Enhancing Resilience Against Jamming Attacks: A Cooperative Anti-Jamming Method Using Direction Estimation

### v1

@@ II System Model
[…]
@@ III-A2 Estimation of Direction of JN’s Channel with Eigenvectors
where $\textbf{u}_{i}$ and $\textbf{v}_{i}$ are the eigenvectors, and $\lambda_{i}$ are the ordered singular values with $\lambda_{1}\geq...\geq\lambda_{K}$. If we assume that the transmitted SNR by the JN, denoted as $\gamma_{t,j}$, is large, this implies that $\textbf{Z}_{t_{p}}\approx\textbf{g}\textbf{a}_{t_{p}}^{T}$ as indicated in equation (8). Consequently, $\textbf{Z}_{t_{p}}$ can be approximated as a rank-one matrix, a technique that has been utilized in various existing studies [37, 38, 39]. This means that it has only one large singular value, and the other singular values tend to zero $\lambda_{i}\rightarrow 0$, where $i=2,...,K$. Thus, using SVD in (11), we have $\textbf{Z}_{t_{p}}\approx\lambda_{1}\textbf{u}_{1}\textbf{v}_{1}^{H}=\overline{\textbf{g}}\|\textbf{g}\|_{2}\textbf{a}_{t_{p}}^{T}$. As a result, we can write $\textbf{u}_{1}\textbf{v}_{1}^{H}-\overline{\textbf{g}}\textbf{a}_{t_{p}}^{T}\|\textbf{g}\|_{2}/\lambda_{1}=\textbf{0}$, where every column should be equal to zero vector, i.e., $\textbf{u}_{1}[\textbf{v}_{1}^{*}]_{n}-\overline{\textbf{g}}[\textbf{a}_{t_{p}}]_{n}\|\textbf{g}\|_{2}/\lambda_{1}=\textbf{0}$, for $n=1,...,N_{t_{p}}-1$. Here, $[\textbf{x}]_{n}$ denotes the $n$-th element of vector x. Accordingly, since $\textbf{v}_{1},\textbf{a}_{t_{p}}\neq\textbf{0}$, we can estimate the direction of g as
@@ III-B Counteracting the Jamming Signal
where we use the least squares (LS) estimation for channel estimation [36]. As the pilot vector is known at the FC, the pseudoinverse of the pilot vector $\textbf{s}_{t_{p}}^{T}$ is $\textbf{s}_{t_{p}}^{+}=\textbf{s}_{t_{p}}^{*}||\textbf{s}_{t_{p}}||^{-2}_{2}$. We multiply (15) by $\textbf{s}_{t_{p}}^{+}$, and using $\textbf{s}_{t_{p}}^{T}\textbf{s}_{t_{p}}^{+}=1$, we have
@@ IV Analytical Discussion on the Proposed Methods
This section analyzes the received SNR from the TN after jamming suppression to evaluate both estimation accuracy and suppression effectiveness. In contrast, Section III used the JN’s SNR to assess jamming direction estimation. The first step is to examine the estimation error of the EV. In the estimation phase, we assume that the JN’s channel direction is a deterministic parameter. The Cramer-Rao lower bound (CRLB) is a useful metric to show the accuracy of the estimation method. To obtain the CRLB, we start with the log-likelihood function (LLF) of the observations in equation (8). Since columns of $\textbf{N}_{t_{p}}$ in (8) follow a zero-mean CSCWG distribution, i.e., $\textbf{n}_{t_{p}}[n]\sim\mathcal{CN}(0,\sigma^{2}\textbf{I})$, the columns of the observation matrix also follow a CSCWG distribution as $\textbf{z}_{t_{p}}[n]\sim\mathcal{CN}(\textbf{g}[\textbf{a}_{t_{p}}]_{n},\sigma^{2}\textbf{I})$. We also employ the notation $[n]$ in $\textbf{z}_{t_{p}}{[n]}$ and $\textbf{n}_{t_{p}}{[n]}$, which are the $n$-th columns of $\textbf{Z}_{t_{p}}$ and $\textbf{N}_{t_{p}}$, to represent their dependence on time. In addition, $[\textbf{a}_{t_{p}}]_{n}$ shows the $n$-th element of the vector $\textbf{a}_{t_{p}}$. The LLF of observations $\textbf{Z}_{t_{p}}$ is
@@
$[\textbf{C}_{\hat{\textbf{g}}}]_{ii}\geq[\tilde{\textbf{I}}(\textbf{g})^{-1}]_{ii},$ (28)
where for the efficient estimator, which has the best performance, the equality holds [41]. In (28), $[\textbf{X}]_{ij}$ shows the element in the $i$-th row and $j$-th column of matrix X. If we approximate the performance of the EV with the performance of the efficient estimator, we can write the covariance of the EV as $\textbf{C}_{\widehat{\overline{\textbf{g}}}}=\tilde{\textbf{I}}(\textbf{g})^{-1}/K=\sigma^{2}_{\epsilon}\textbf{I}$, where $\sigma^{2}_{\epsilon}=\sigma^{2}/K\sum_{n=1}^{N_{t_{p}}-1}|\textbf{j}_{t_{p}}^{H}{\boldsymbol{\psi}_{n}}|^{2}$, and for simplicity, we assumed that $\|\textbf{g}\|_{2}=\sqrt{K}$ and $\overline{\textbf{g}}=\textbf{g}/\sqrt{K}$. In addition, since the EV method is an unbiased estimator that is a linear function of the observation matrix $\textbf{Z}_{t_{p}}$, we have $\widehat{\overline{\textbf{g}}}\sim\mathcal{CN}(\overline{\textbf{g}},\sigma^{2}_{\epsilon}\textbf{I})$ [42]. Consequently, the estimation error for the EV method is $\textbf{e}=\widehat{\overline{\textbf{g}}}-\overline{\textbf{g}}$, where $\textbf{e}\sim\mathcal{CN}(\textbf{0},\sigma^{2}_{\epsilon}\textbf{I})$. We can now integrate the error in the estimation of the JN’s channel direction during the jamming removal phase. In the best-case scenario, the desirable $\textbf{G}^{\perp}$ satisfies ${\textbf{G}^{\perp}}^{H}\overline{\textbf{g}}=\textbf{0}$. However, since we use the estimation of the JN’s channel direction, we have to consider the estimation error. Therefore, we can write ${\textbf{G}^{\perp}}^{H}\widehat{\overline{\textbf{g}}}={\textbf{G}^{\perp}}^{H}\overline{\textbf{g}}+{\textbf{G}^{\perp}}^{H}\textbf{e}=\textbf{0}$, and consequently, it can be concluded that
@@
Evidently, as columns of $\textbf{G}^{\perp}$ are normalized, the transformed vector of estimation error also follows $-{\textbf{G}^{\perp}}^{H}\textbf{e}\sim\mathcal{CN}(\textbf{0},\sigma^{2}_{\epsilon}\textbf{I})$, while its size is $K-1$. Using (2) and its transformed version with ${\textbf{G}^{\perp}}$, we have
@@
This analytical formula for the received SNR after the removal of the JN’s signal, enables us to calculate the outage probability as $P_{out}=\mathrm{Pr}\{\gamma_{r,s}<\gamma_{th}\}$. By assuming that the channel vector between the TN and SNs follows a Rayleigh fading model and a normal distribution with $\textbf{h}\sim\mathcal{CN}(\textbf{0},\textbf{I})$, we can argue that $\|\textbf{f}\|_{2}^{2}$ follows a chi-squared distribution with $2(K-1)$ real DoF. Denoting the cumulative distribution function (CDF) of a real-valued random variable with chi-squared distribution and with $k$ real DoF by $\chi^{2}_{k}(.)$, the outage probability for the EV-CAJ is
@@ VI Simulation Results
Threshold SNR for Outage $\gamma_{th}$10 dB
Symbol Duration $t_{s}$
@@
JN’s Channel Vector Coherence Time $t_{c,g}$
TN’s Channel Vector Coherence Time $t_{c,h}$
@@ VI-A Estimation Performance
where ||$\cdot$|| ⋅ | is the element-wise absolute value of the vector. We use the absolute value because we are specifically interested in the energy difference in each dimension of the vectors, which acts as a metric for the error in the estimation of the direction.
@@ VI-C Symbol Error Probability
Fig. 9 shows the simulated $P_{se}$ versus $\gamma_{t,s}$ for $K=4$ and $N_{t_{p}}=50$. We observe that the proposed EV-CAJ attains $P_{es}=10^{-3}$ with $\gamma_{t,s}$ roughly $2.5$ dB higher than the No JN case. The NLS-CAJ requires an SNR about $7$ dB larger than the case where no JN is present to attain the same $P_{es}$. It is also observed that methods based on ZFIC show inferior performances even though perfect CSI of the JN’s channel is used. This figure also shows that the EV method achieves a very accurate estimation of the jamming channel since the performance of the EV-CAJ is very close to that of the CAJ with perfect CSI on the JN’s channel. In addition, the curve of the EV-CAJ method with $K=4$ is very similar to that of the No JN case with $K=3$. The reason behind this is that we removed the jamming with the use of $\textbf{G}^{\perp}$, which transforms the received signal into a new space with $K-1$ dimensions. This new space is the orthogonal complement space of the line space spanned by g. Therefore, we can effectively eliminate a large portion of the jamming signal at the cost of losing one DoF.
@@ VI-E CAJ Under Time-Varying Channels
To further investigate the joint effect of $\gamma_{t,j}$ and $\tau_{g}$, Figs. 16 and 17 are presented. These two figures contain contours of $P_{se}$ attained by the EV-CAJ method under various $\tau_{g}$ and different transmit SNRs by the JN and the TN. Both figures include five subfigures where each subfigure presents a $\textbf{g}[n]$ with a specific $\tau_{g}$ from $\tau_{g}\in\{0,0.01,0.04,0.1$, 0.5}0.5\}0.5 }. Color bars in these figures show the logarithm of $P_{se}$. For example, the blue color in this color bar shows the number $-3$ which represents the $P_{se}=10^{-3}$. Thus, the blue area in each subfigure is the desirable area since it offers a small error rate. In Fig. 16, for $K=4$, the blue area is positioned in a higher $\gamma_{t,s}$ compared to Fig. 17, with $K=16$. Interestingly, for the TI channel where $\tau_{g}=0$, the transmit SNR by the JN has no impact on the performance of the EV-CAJ method, as shown in subfigure (a). However, by increasing $\tau_{g}$ and the speed of variation in $\textbf{g}[n]$, we can see that areas in the top right of the sub-figures gradually turn into red. This implies that if the JN’s signal is very strong, the channel variation is further damaging. In contrast, a signal transmitted by the JN with a weaker SNR can be removed, even under a fast-fading channel. This is why the blue area for $\gamma_{t,j}\leq 20$ dB is quite constant for $\tau_{g}\leq 0.04$. We can conclude that the EV-CAJ method can also effectively combat JNs with TV channels depending on the power of the JN’s signal and its coherence time. Based on these results, we face a trade-off in the selection of the number of pilots and data symbols. If the JN’s signal is very strong, and $\tau_{g}$ is large, the EV-CAJ can handle the jamming attack effectively by decreasing $t_{d}$ and transmitting pilot blocks more frequently. This approach reduces $\tau_{g}$, which leads to a significantly smaller $P_{se}$ under a strong jamming signal at the expense of a lower data rate. On the other hand, if $\gamma_{t,j}$ and $\tau_{g}$ are small (a similar scenario is depicted at the bottom of subfigures (a) and (b)), the probability of error remains still low even with fewer numbers of pilot symbols, allowing us to benefit from the increased $t_{d}$ as $t_{p}$ becomes smaller. Accordingly, we observe that for $\tau_{g}<0.01$, EV-CAJ performs effectively even in the presence of strong jamming signals. Therefore, a general guideline for selecting the duration of the data block could be $t_{d}<0.01t_{c,g}$. If the jamming signal is moderate ($\gamma_{t,j}<15$ dB), this constraint can be further relaxed to $t_{d}<0.04t_{c,g}$, allowing for a longer data block.

### v2

@@ II System Model
where $\textbf{h}=[h_{1},...,h_{k}]^{T}$, $\textbf{g}=[g_{1},...,g_{k}]^{T}$, $\textbf{W}_{t}=[\textbf{w}_{t,1},...,\textbf{w}_{t,K}]^{T}$, and $\textbf{W}_{t}\in\mathbb{C}^{K\times N_{t}}$ . The vector h includes the channel gains between SNs and the TN, and g denotes the channel gains vector between SNs and the JN. Additionally, we assume that the channel vectors and sequence of signals are not identical, i.e., $\textbf{h}\neq\textbf{g}$, and $\textbf{s}\neq\textbf{j}$, which implies that $\mathrm{rank}(\textbf{h}\textbf{s}_{t}^{T}+\textbf{g}\textbf{j}_{t}^{T})=2$. Moreover, $\|\textbf{s}_{t}\|_{2}=\sqrt{P_{t,s}},\|\textbf{j}_{t}\|_{2}=\sqrt{P_{t,j}}$, where $\|\cdot\|_{2}$ is the Euclidean norm. For each column of the noise matrix, we assume that the noise is circularly symmetric complex white Gaussian (CSCWG), and we have $\textbf{w}_{t}[n]\sim\mathcal{CN}(0,\sigma^{2}\textbf{I})$, where $\textbf{w}_{t}[n]$ is the $n$-th column of $\textbf{W}_{t}$, and shows the values of noise in all SNs at the $n$-th moment.
@@ III-A2 Estimation of Direction of JN’s Channel with Eigenvectors
where $\textbf{u}_{i}$ and $\textbf{v}_{i}$ are the eigenvectors, and $\lambda_{i}$ are the ordered singular values with $\lambda_{1}\geq...\geq\lambda_{K}$. If we assume that the transmitted SNR by the JN, denoted as $\gamma_{t,j}$, is large, this implies that $\textbf{Z}_{t_{p}}\approx\textbf{g}\textbf{a}_{t_{p}}^{T}$ as indicated in equation (8). Consequently, $\textbf{Z}_{t_{p}}$ can be approximated as a rank-one matrix, a technique that has been utilized in various existing studies [37, 38, 39]. This means that it has only one large singular value, and the other singular values tend to zero $\lambda_{i}\rightarrow 0$, where i=2,,Ki=2,...,K. Thus, using SVD in (11), we have $\textbf{Z}_{t_{p}}\approx\lambda_{1}\textbf{u}_{1}\textbf{v}_{1}^{H}=\overline{\textbf{g}}\|\textbf{g}\|_{2}\textbf{a}_{t_{p}}^{T}$. As a result, we can write $\textbf{u}_{1}\textbf{v}_{1}^{H}-\overline{\textbf{g}}\textbf{a}_{t_{p}}^{T}\|\textbf{g}\|_{2}/\lambda_{1}=\textbf{0}$, where every column should be equal to zero vector, i.e., $\textbf{u}_{1}[\textbf{v}_{1}^{*}]_{n}-\overline{\textbf{g}}[\textbf{a}_{t_{p}}]_{n}\|\textbf{g}\|_{2}/\lambda_{1}=\textbf{0}$, for $n=1,...,N_{t_{p}}-1$. Here, $\textbf{x}]_{n}$ denotes the $n$-th element of vector x. Accordingly, since $\textbf{v}_{1},\textbf{a}_{t_{p}}\neq\textbf{0}$, we can estimate the direction of g as
@@ III-B Counteracting the Jamming Signal
[…]
@@ IV Analytical Discussion on the Proposed Methods
This section analyzes the received SNR from the TN after jamming suppression to evaluate both estimation accuracy and suppression effectiveness. In contrast, Section III used the JN’s SNR to assess jamming direction estimation. The first step is to examine the estimation error of the EV. In the estimation phase, we assume that the JN’s channel direction is a deterministic parameter. The Cramer-Rao lower bound (CRLB) is a useful metric to show the accuracy of the estimation method. To obtain the CRLB, we start with the log-likelihood function (LLF) of the observations in equation (8). Since columns of $\textbf{N}_{t_{p}}$ in (8) follow a zero-mean CSCWG distribution, i.e., $\textbf{n}_{t_{p}}[n]\sim\mathcal{CN}(0,\sigma^{2}\textbf{I})$, the columns of the observation matrix also follow a CSCWG distribution as $\textbf{z}_{t_{p}}[n]\sim\mathcal{CN}(\textbf{g}[\textbf{a}_{t_{p}}]_{n},\sigma^{2}\textbf{I})$. We also employ the notation $[n]$ in $\textbf{z}_{t_{p}}{[n]}$ and $\textbf{n}_{t_{p}}{[n]}$, which are the $n$-th columns of $\textbf{Z}_{t_{p}}$ and $\textbf{N}_{t_{p}}$, to represent their dependence on time. In addition, $\textbf{a}_{t_{p}}]_{n}$ shows the $n$-th element of the vector $\textbf{a}_{t_{p}}$. The LLF of observations $\textbf{Z}_{t_{p}}$ is
@@
$\textbf{C}_{\hat{\textbf{g}}}]_{ii}\geq[\tilde{\textbf{I}}(\textbf{g})^{-1}]_{ii},$ (28)
where for the efficient estimator, which has the best performance, the equality holds [41]. In (28), $\textbf{X}]_{ij}$ shows the element in the $i$-th row and $j$-th column of matrix X. If we approximate the performance of the EV with the performance of the efficient estimator, we can write the covariance of the EV as $\textbf{C}_{\widehat{\overline{\textbf{g}}}}=\tilde{\textbf{I}}(\textbf{g})^{-1}/K=\sigma^{2}_{\epsilon}\textbf{I}$, where $\sigma^{2}_{\epsilon}=\sigma^{2}/K\sum_{n=1}^{N_{t_{p}}-1}|\textbf{j}_{t_{p}}^{H}{\boldsymbol{\psi}_{n}}|^{2}$, and for simplicity, we assumed that $\|\textbf{g}\|_{2}=\sqrt{K}$ and $\overline{\textbf{g}}=\textbf{g}/\sqrt{K}$. In addition, since the EV method is an unbiased estimator that is a linear function of the observation matrix $\textbf{Z}_{t_{p}}$, we have $\widehat{\overline{\textbf{g}}}\sim\mathcal{CN}(\overline{\textbf{g}},\sigma^{2}_{\epsilon}\textbf{I})$ [42]. Consequently, the estimation error for the EV method is $\textbf{e}=\widehat{\overline{\textbf{g}}}-\overline{\textbf{g}}$, where $\textbf{e}\sim\mathcal{CN}(\textbf{0},\sigma^{2}_{\epsilon}\textbf{I})$. We can now integrate the error in the estimation of the JN’s channel direction during the jamming removal phase. In the best-case scenario, the desirable $\textbf{G}^{\perp}$ satisfies ${\textbf{G}^{\perp}}^{H}\overline{\textbf{g}}=\textbf{0}$. However, since we use the estimation of the JN’s channel direction, we have to consider the estimation error. Therefore, we can write ${\textbf{G}^{\perp}}^{H}\widehat{\overline{\textbf{g}}}={\textbf{G}^{\perp}}^{H}\overline{\textbf{g}}+{\textbf{G}^{\perp}}^{H}\textbf{e}=\textbf{0}$, and consequently, it can be concluded that
@@
Evidently, as columns of $\textbf{G}^{\perp}$ are normalized, the transformed vector of estimation error also follows ${\textbf{G}^{\perp}}^{H}\textbf{e}\sim\mathcal{CN}(\textbf{0},\sigma^{2}_{\epsilon}\textbf{I})$, while its size is K1K-1. Using (2) and its transformed version with $\textbf{G}^{\perp}$}, we have
@@
This analytical formula for the received SNR after the removal of the JN’s signal, enables us to calculate the outage probability as $P_{out}=\mathrm{Pr}\{\gamma_{r,s}<\gamma_{th}\}$. By assuming that the channel vector between the TN and SNs follows a Rayleigh fading model and a normal distribution with $\textbf{h}\sim\mathcal{CN}(\textbf{0},\textbf{I})$, we can argue that $\|\textbf{f}\|_{2}^{2}$ follows a chi-squared distribution with 2(K1)2(K-1) real DoF. Denoting the cumulative distribution function (CDF) of a real-valued random variable with chi-squared distribution and with $k$ real DoF by $\chi^{2}_{k}(.)$, the outage probability for the EV-CAJ is
@@ VI Simulation Results
Threshold SNR for Outage $\gamma_{th}$ -10 dB
Symbol Duration $t_{s}$ -
@@
JN’s Channel Vector Coherence Time $t_{c,g}$ -
TN’s Channel Vector Coherence Time $t_{c,h}$ -
@@ VI-A Estimation Performance
where $|\cdot|$ is the element-wise absolute value of the vector. We use the absolute value because we are specifically interested in the energy difference in each dimension of the vectors, which acts as a metric for the error in the estimation of the direction.
@@ VI-C Symbol Error Probability
Fig. 9 shows the simulated $P_{se}$ versus $\gamma_{t,s}$ for $K=4$ and $N_{t_{p}}=50$. We observe that the proposed EV-CAJ attains $P_{es}=10^{-3}$ with $\gamma_{t,s}$ roughly $2.5$ dB higher than the No JN case. The NLS-CAJ requires an SNR about $7$ dB larger than the case where no JN is present to attain the same $P_{es}$. It is also observed that methods based on ZFIC show inferior performances even though perfect CSI of the JN’s channel is used. This figure also shows that the EV method achieves a very accurate estimation of the jamming channel since the performance of the EV-CAJ is very close to that of the CAJ with perfect CSI on the JN’s channel. In addition, the curve of the EV-CAJ method with $K=4$ is very similar to that of the No JN case with $K=3$. The reason behind this is that we removed the jamming with the use of $\textbf{G}^{\perp}$, which transforms the received signal into a new space with K1K-1 dimensions. This new space is the orthogonal complement space of the line space spanned by g. Therefore, we can effectively eliminate a large portion of the jamming signal at the cost of losing one DoF.
@@ VI-E CAJ Under Time-Varying Channels
To further investigate the joint effect of $\gamma_{t,j}$ and $\tau_{g}$, Figs. 16 and 17 are presented. These two figures contain contours of $P_{se}$ attained by the EV-CAJ method under various $\tau_{g}$ and different transmit SNRs by the JN and the TN. Both figures include five subfigures where each subfigure presents a $\textbf{g}[n]$ with a specific $\tau_{g}$ from $\tau_{g}\in\{0,0.01,0.04,0.1$, 0.5}0.5\}. Color bars in these figures show the logarithm of $P_{se}$. For example, the blue color in this color bar shows the number 3-3 which represents the $P_{se}=10^{-3}$. Thus, the blue area in each subfigure is the desirable area since it offers a small error rate. In Fig. 16, for $K=4$, the blue area is positioned in a higher $\gamma_{t,s}$ compared to Fig. 17, with $K=16$. Interestingly, for the TI channel where $\tau_{g}=0$, the transmit SNR by the JN has no impact on the performance of the EV-CAJ method, as shown in subfigure (a). However, by increasing $\tau_{g}$ and the speed of variation in $\textbf{g}[n]$, we can see that areas in the top right of the sub-figures gradually turn into red. This implies that if the JN’s signal is very strong, the channel variation is further damaging. In contrast, a signal transmitted by the JN with a weaker SNR can be removed, even under a fast-fading channel. This is why the blue area for $\gamma_{t,j}\leq 20$ dB is quite constant for $\tau_{g}\leq 0.04$. We can conclude that the EV-CAJ method can also effectively combat JNs with TV channels depending on the power of the JN’s signal and its coherence time. Based on these results, we face a trade-off in the selection of the number of pilots and data symbols. If the JN’s signal is very strong, and $\tau_{g}$ is large, the EV-CAJ can handle the jamming attack effectively by decreasing $t_{d}$ and transmitting pilot blocks more frequently. This approach reduces $\tau_{g}$, which leads to a significantly smaller $P_{se}$ under a strong jamming signal at the expense of a lower data rate. On the other hand, if $\gamma_{t,j}$ and $\tau_{g}$ are small (a similar scenario is depicted at the bottom of subfigures (a) and (b)), the probability of error remains still low even with fewer numbers of pilot symbols, allowing us to benefit from the increased $t_{d}$ as $t_{p}$ becomes smaller. Accordingly, we observe that for $\tau_{g}<0.01$, EV-CAJ performs effectively even in the presence of strong jamming signals. Therefore, a general guideline for selecting the duration of the data block could be $t_{d}<0.01t_{c,g}$. If the jamming signal is moderate ($\gamma_{t,j}<15$ dB), this constraint can be further relaxed to $t_{d}<0.04t_{c,g}$, allowing for a longer data block.